Write in slope intercept form the equation of the line through (-1,2) perpendicular to 2x-4y=6
Answer by robertb(5830)

The unknown line must have the form 4x + 2y = c, where c is an undetermined coefficient.
To find c, use the coordinates of the given point:
4*-1 + 2*2 = -4 + 4 = 0.
Hence, the equation is 4x + 2y = 0, or y = -2x + 0, in slope-intercept form.